# Inverse Garment and Pattern Modeling with a Differentiable Simulator

### 11 Mar 2024 v1

@@ 7 Matrix Assembly Acceleration
8 Ablation Study
In order to examine the contributions of individual elements in our proposed method, we conducted an ablation study (Table 3), assessing their impact on the overall performance. Our baseline model achieved a Chamfer distance precision (CF) of 0.1103 (L2 mm). By replacing the curvature-weighted Chamfer loss with the vanilla Chamfer, we obtained an increase of CF to 0.115. When we remove the seam consistency term from the loss, CF increases to 0.117. It affirms the importance of both loss terms in capturing the geometry. When the optimization of physical parameters was disabled, CF has increased to 0.113, confirming the importance of incorporating physical properties in replicating complex garment shapes.
Fig. 13 illustrates the reconstructed models obtained from the ablation study. We observe that our model bears the closest resemblance to the target. The use of weighted Chamfer distance allows for better capture of the armpit region and the lower part of the sleeve (Fig. 13(e) vs. Fig. 13(b). The absence of seam loss causes ”puckered seam” around the shoulders, induced by the extra tension on the panel with shorter seam length (Fig. 13(c) vs. Fig. 13(e). The optimization of physical parameters helps to recover fine wrinkles, as well as more plausible pattern estimation. As we can see in Fig. 13(d), the larger hem-width of the dress has been solely attributed to the larger width of the pattern when the material parameters were disregarded during the optimization.
Table 3: Alblation Study.
Method CF NOR
@@ 8 Ablation Study
Ours(w/o seaming loss) 0.117 0.076
@@
Ours 0.1103 0.075
Figure 13: Results from our ablation study on loss terms. (a) Target; (b) Chamfer loss without curvature weights (with vanilla Chamfer); (c) without seam loss (d) without physical parameters (e) baseline.
9 Limitations
Our approach presents several limitations that suggest avenues for future exploration. First, the iterative optimization process involving forward and reverse simulation is time-consuming. Acceleration methods can be employed to guide faster convergence in optimization processes [23]. Second, although our linear grading scheme effectively adjusts the base model to align with the target, our results can be considerably influenced by initial values for the optimization, as well as predefined design choices such as mesh resolution and the identification of control points on the panels. Combining our model with a learning-based method to achieve good initialization would be an interesting future investigation. Such hybrid approaches can also be adopted to address other issues like special design choices, which is currently supported through manual customization of base models.
10 Pattern Symmetry Detection
A detailed algorithm description of inter-panel and intra-panel symmetry detection can be found in Algorithm 1.
11 Linear Grading
Once an initial draped shape of the 3D base garment has been simulated, the 2D pattern is first adjusted by taking into account the difference in circumferences and their axial distances along the bone between corresponding open contours measured in 3D. Contours around the hip or waist, cuff contours on the arm, are typically concerned by this linear grading step. In Fig. 14, we illustrate the editing of 2D panel curves by using the measurements of cuff contours in 3D. The distance between the target cuff $O_{t}$ and the source $O_{s}$ measured along the arm bones, together with their difference in circumference, determines the amount of displacement of control points on the corresponding open curves in the pattern. In a 2D panel, the direction of length change $\vec{d}$ is derived by computing and normalizing the midpoint of the two control points along the open curve minus the average of the remaining control points, while the width-changing vectors are computed by subtracting each control point from the other and normalizing it. Readers can refer to Algorithm 2 for a detailed procedure for the distance measurement along the skeleton.
Figure 14: An example of linear grading. The axial distance between open contours from the source mesh (left) to the target mesh (right) is used to displace the corresponding open curve in the corresponding panel.
12 More Experimental Results
Fig. 15 shows additional results we obtained on different styles of pants. It clearly shows that our method can reconstruct both tight and loose clothes.
Figure 15: Our method can handle different garment styles: (a) leggings; and (b) baggy pants. For each style, our reconstructed model (top left), the ground-truth garment (top right), and the estimated pattern (bottom) are illustrated.
[…]
Algorithm 1 Symmetry Detection.
[…]
Algorithm 2 Axial Distances of Open Contours.

### 15 Mar 2024 v2

@@ Appendix 0.B Matrix assembly acceleration
Appendix 0.C Ablation study
Here we report the results of our ablation study, where we examine the contributions of individual components to the overall performance (Table 3). Our baseline model achieved a Chamfer distance precision (CF) of 0.1103 (in mm) and a reduced cosine distance of normals (NOR) of 0.075. It outperforms other settings where the curvature-weighted Chamfer loss is replaced with the vanilla Chamfer (CF: 0.115, NOR:0.085), when the seam consistency loss term is removed (CF:0.117, NOR:0.076), or when the optimization of physical parameters was disabled (CF: 0.113, NOR: 0.081). These results confirm the importance of both loss terms and the integration of physical properties in the optimization process.
Fig. 10 illustrates the reconstructed models obtained from the ablation study. We observe that our model (Fig. 10(e)) bears the closest visual resemblance to the target. The use of weighted Chamfer distance allows for better capture of the armpit region and the lower part of the sleeve (Fig. 10(b)). The absence of seam loss leads to a puckered seam around the shoulder, resulting from the extra tension exerted on the the shorter seam (Fig. 10(c)). The optimization of physical parameters helps to recover fine wrinkles, as well as more plausible pattern estimation. As we can see in Fig. 10(d), the “A-line” shape of the body/dress has been solely attributed to the increasing panel width along the torso, when the material parameters were disregarded during the optimization.
Table 3: Ablation Study.
Method Chamfer distance (CF) Normal difference (NOR)
@@ Appendix 0.C Ablation study
Ours(w/o seam loss) 0.117 0.076
@@
Ours 0.110 0.075
Figure 10: Results of our ablation study on loss terms. (a) Target; (b) Chamfer loss without curvature weights (i.e. vanilla Chamfer); (c) without seam loss; (d) without physical parameters; (e) Ours.
Appendix 0.D Pattern symmetry detection
A detailed algorithmic description of inter-panel and intra-panel symmetry detection can be found in Algorithm 1.
Appendix 0.E Linear grading
Once the initial draping of the 3D base garment has been simulated, the 2D pattern is adjusted to minimize the difference in both circumferences and axial distances along the bone between corresponding open contours in 3D. Contours around the hip or waist, cuff contours on the arm, are typically used in this linear grading step. In Fig. 11, the distance between the target cuff $O_{t}$ and the source $O_{s}$ measured along the arm bones, together with the difference in their circumferences, determines the amount of the displacement $|\overrightarrow{\vec{d}}|$ of control points on the corresponding open curve (in blue) of the pattern. In the 2D panel, the direction of $\overrightarrow{\vec{d}}$ is derived by computing and normalizing the midpoint of the two control points on the open curve minus the average of the remaining control points, while the width-changing vectors are computed by subtracting each control point from the other and normalizing it. Readers can refer to Algorithm 2 for a detailed procedure on pairing and computing axial distances for open contours.
Figure 11: An example of linear grading. The axial distance $\overrightarrow{\vec{d}}$ between open contours from the source mesh (left) to the target mesh (right) is used to displace the corresponding open curve (in blue) in the 2D panel.
Appendix 0.F Limitations
Our approach presents several limitations that suggest avenues for future exploration. First, the iterative optimization process involving forward and reverse simulation is time-consuming. Further acceleration can be employed to achieve faster convergence in optimization processes [23]. Second, although our linear grading scheme effectively adjusts the base model to align with the target prior to the optimization, our final results can be sensitive to initial values, potentially resulting in different local minima, as well as to predefined design choices such as mesh resolution and the identification of control points on the panels. Combining our model with a learning-based method to achieve good initialization would be an interesting future investigation.